OpenStudy (anonymous):

ABCD is a parallelogram, lettered anticlockwise, such that A and Care the points (−1, 5) and (5, 1) respectively. a) Find the coordinates of the midpoint of AC. b) Given that BD is parallel to the line whose equation is y+5x=2, find the equation of BD. c) Given that BC is perpendicular to AC, find: i. the equation of BC ii. the coordinates of B iii. the coordinates of D.

OpenStudy (anonymous):

im not sure what to do with c.iii. finding the coor of D

OpenStudy (anonymous):

a)(2,3) b)y-3=-5(x-2) y=-5x+13

OpenStudy (anonymous):

yeah for c?

OpenStudy (anonymous):

C)you know the slope for AC right??slope of BC=-1/slope of AC you know point C(x1,y1) too....so y-y1=slope of BC(x-x1)

OpenStudy (anonymous):

yteah

OpenStudy (anonymous):

i just cant do iii. thats it

OpenStudy (anonymous):

slope of BC is same as AD right??(since its a parallelogram)...you know point A and slope of AD...so find D...

OpenStudy (anonymous):

got it??

OpenStudy (anonymous):

no sorry then what i do aftert?

OpenStudy (sirm3d):

@powerangers69 use the fact that the midpoint (2,3) in problem (a) is also the midpoint of BD.

OpenStudy (anonymous):

hmmmm